### v2

@@ V Numerical Results and Discussion
Fig. 5(a) shows the beampattern gains of the JPT scheme with varying $\rho$. When $\rho=1$, the mainlobe gain is minimum. This is because the optimization objective in the JPT scheme reduces to minimizing the gains at $D$’s direction, which only maximizes the eavesdropping rate. To characterize the effect of $\rho$ on the sensing performance in the high-$\rho$ region, we increase the density of the points when $\rho>0.9$ in Fig. 5. As $\rho$ decreases, the mainlobe gain increases and sensing performance is enhanced, but the increase is limited. This is because both PE and sensing performance are maximized. Figs. 5(b)-5(d) plot the determinant of CRB with varying $\rho$, respectively. One interesting conclusion can be observed that, as $\rho$ increases, the CRB keeps constant until it suddenly increases in the high-$\rho$ region. This is because the CRB is of a small order of magnitude and is less sensitive to changes in the covariance matrix $\mathbf{R}$ and a higher $\rho$ denotes the PE will be the main part of the objective function in the JPT scheme. Comparing Figs. 5(b)-5(d), one can find that the sensing performance for all the schemes is improved as $P_{0}$ increases, which also can be found in Fig. 2(b). Furthermore, the JPT’s sensing performance with higher power ($P_{0}=30$ dBm) outperforms that of the TSC. This is because, in the JPT scheme, the beampatterns gain in $D$’s direction is kept constant to maximize the eavesdropping rate. Then, more power is allocated to the mainlobe, resulting in better sensing performance.
(a) The eavesdropping rate with $P_{0}=26$ dBm.
(b) The eavesdropping rate with $P_{0}=28$ dBm.
(c) The eavesdropping rate with $P_{0}=30$ dBm.
@@
Figs. 6(a)-6(c) plot the eavesdropping rate with varying $\rho$, respectively. We can observe that, in the high-$\rho$ region, the eavesdropping rate of the JPT increases to the maximum, which equals that of PEC. Moreover, with the increase of $P_{0}$, the minimum $\rho$ to achieve the maximum eavesdropping rate decreases. This is because a large $\rho$ implies a greater preference for PE. For TSC, the eavesdropping rate decreases with the increase of $P_{0}$, which also can be found in Fig. 2(c). It must be noted that, in Figs. 5 and 6, $\rho\neq 0$ denotes that the JPT scheme enables both PE and TS simultaneously.
@@
Figs. 7(a) plots the beampattern gains with uncertainty angles for the JPT scheme. We can find that, with the increase of ${\Delta\theta}$, the mainlobe width becomes more expansive and the gain decreases because the energy is spread over a larger area. However, in the $D$’s direction, the gain must be not less than the requirement. Fig. 7(b) plots the beampattern gains when the interference power is equal to zero. Because the quality of the eavesdropping channel is better than that of the illegal communication channel, it is not necessary to transmit power in $D$’s direction.
@@ VI Conclusion
This work studied the beamforming design in the joint PE and TS system wherein the multi-antenna BS transmits the signal to realize TS, which was utilized as AN to achieve PE successful. Firstly, two waveform design schemes were proposed to maximize the eavesdropping rate or minimize the CRB only. Then, three waveform design schemes were investigated by formulating PE-centric, TS-centric, and normalized weighted optimization problems. By omitting the rank-one constraint, the non-convex problem was solved by the SDR technique. The SROCR method was utilized to solve continuously and iteratively to obtain the rank-one beamforming solution. In this work, it is assumed that the CSI and location of the illegal communication links were available at the eavesdropper. In some scenarios, the BS needs to sense and estimate the CSI and the location of the illegal communication transmitter and receivers, which will be part of future work.

### v3

@@ V Numerical Results and Discussion
Fig. 5(a) shows the beampattern gains of the JPT scheme with varying $\rho$. When $\rho=1$, the mainlobe gain is minimum because the optimization objective in the JPT scheme reduces to minimizing the gains at $D$’s direction, which only maximizes the ER. Figs. 5(b)-5(d) plot the DCRB with varying $\rho$, respectively. One interesting conclusion can be observed that, in the JPT scheme, the DCRB remains constant for all $\rho$ values except at the points where $\rho=0$ and $\rho=1$. The reason is that when $\rho=0$, the JPT scheme degenerates to the TSC scheme; when $\rho=1$, it degenerates to the PEC scheme. Comparing Figs. 5(b)-5(d), one can find that the sensing performance for all the schemes is improved as $P_{0}$ increases, which can also be found in Fig. 2(b). Furthermore, the JPT’s sensing performance with $\rho$ outperforms that of the TSC. This is because, in the JPT scheme, the beampattern gain in $D$’s direction is kept constant to maximize the ER. Then, more power is allocated to the mainlobe, resulting in better sensing performance.
(a) The ER with $P_{0}=28$ dBm.
(b) The ER with $P_{0}=30$ dBm.
(c) The ER with $P_{0}=32$ dBm.
@@
Figs. 6(a)-6(c) plot the ER with varying $\rho$, respectively. It can be observed that, when $\rho\neq 0$, the ER of JPT is equal to that of PEC. Compared among Figs. 6(a)-6(c), the ER of all the schemes exhibits a relatively weak dependence on the $P_{0}$, which can also be found in Fig. 2(c). It should be noted that the ER is equal to $R_{D}$, which is given in (11). Maxmimzing ER is equal to maxmimzing $R_{D}$, which is expressed as ${\gamma_{D}}=\frac{{{P_{S}}{{\left|{{h_{SD}}}\right|}^{2}}}}{{\sigma_{D}^{2}}}$ and is independent of $P_{0}$. The ER of TSC shown in Figs. 6(a)-6(c) has a certain degree of randomness because the TSC scheme does not take into account maximizing ER, and ER is considered as a constraint (21f). It must be noted that the results in Figs. 5 and 6 clearly demonstrate that there is no trade-off between PE and TS. The core reason is as follows: the proposed joint optimization framework multiplexes radar waveforms to simultaneously transmit sensing signals and AN, which not only enables TS but also serves as AN to achieve successful PE. Specifically, by constraining the beam gain in the $D$’s direction, the ER can be maximized while ensuring successful PE. Meanwhile, the constrained energy in this direction enhances the degrees of freedom in energy distribution, thereby improving the TS performance. Consequently, the JPT design outperforms that of TSC and PEC.
@@
Fig. 7(a) plots the beampattern gains with uncertain angles for the JPT scheme. We can find that, with the increase of $\Delta\theta$}, the mainlobe width becomes more expansive and the gain decreases because the energy is spread over a larger area. However, in the $D$’s direction, the gain must be not less than the requirement. Fig. 7(b) plots the beampattern gains when the interference power is equal to zero. Because the quality of the eavesdropping channel is better than that of the illegal communication channel, it is not necessary to transmit power in $D$’s direction.
@@ VI Conclusion
This work studied the beamforming design in the joint PE and TS system wherein the multi-antenna BS transmits the signal to realize TS, which was utilized as AN to achieve PE successfully. Firstly, two waveform design schemes were proposed to maximize the ER or minimize the DCRB only. Then, three waveform design schemes were investigated by formulating PE-centric, TS-centric, and normalized weighted optimization problems. By omitting the rank-one constraint, the non-convex problem was solved by the SDR technique. The SROCR method was utilized to solve continuously and iteratively to obtain the rank-one beamforming solution. In this work, it is assumed that the CSI and location of the illegal communication links are available. However, the channel estimation errors for illicit eavesdroppers also significantly affect practical results. The scenarios wherein the BS senses the illegal transmitters and receivers, and/or with the imperfect CSI of the unlawful links, will be part of future work.
Appendix A Derivation of CRB
The Fisher information matrix for ${\bm{\zeta}}=\left[{{{\bm{\theta}}},{{\bm{\beta}}_{R}},{{{\bm{\beta}}}_{I}}}\right]$ is expressed as
${\textbf{F}}=\left[{{\textbf{F}}{{\left({\bm{\zeta}}\right)}_{ij}}}\right]$
$=\left[{\frac{2}{{{\sigma^{2}}}}{\textrm{Re}{\mathrm{tr}}}\left[{\frac{{\partial{{\left({{{\textbf{A}}_{r}}\left({\bm{\theta}}\right){\bm{\beta}}{\textbf{A}}_{t}^{H}\left(\theta\right){{\textbf{x}}_{E}}}\right)}^{H}}}}{{\partial{\zeta_{i}}}}\frac{{\partial\left({{{\textbf{A}}_{r}}\left({\bm{\theta}}\right){\bm{\beta}}{\textbf{A}}_{t}^{H}\left({\bm{\theta}}\right){{\textbf{x}}_{E}}}\right)}}{{\partial{\zeta_{j}}}}}\right]}\right]$
[…]
where ${\mathbf{F}}{\left({\bm{\xi}}\right)_{{{ij}}}}={\mathrm{tr}}\left[{{{\mathbf{C}}^{-1}}\left({\bm{\xi}}\right)\frac{{\partial{\mathbf{C}}\left({\bm{\xi}}\right)}}{{\partial{\xi_{i}}}}{{\mathbf{C}}^{-1}}\left({\bm{\xi}}\right)\frac{{\partial{\mathbf{C}}\left({\bm{\xi}}\right)}}{{\partial{\xi_{j}}}}}\right]+2{\mathop{\mathrm{Re}}\nolimits}\left[{\frac{{\partial{{\bm{\mu}}^{H}}\left({\bm{\xi}}\right)}}{{\partial{\xi_{i}}}}{{\mathbf{C}}^{-1}}\left({\bm{\xi}}\right)\frac{{\partial{\bm{\mu}}\left({\bm{\xi}}\right)}}{{\partial{\xi_{j}}}}}\right]$ is a KKK*K matrix.
1) The first element is expressed as
[…]
where $F\left({{\theta_{i}},{\theta_{j}}}\right)$ is obtained as (27), shown at the top of the next page,
$F\left({{\theta_{i}},{\theta_{j}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}{\mathop{\mathrm{tr}}\nolimits}\left[{\frac{{\partial{{\left({{{\mathbf{A}}_{r}}{\bm{\beta A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}^{H}}}}{{\partial{\theta_{i}}}}\frac{{\partial\left({{{\mathbf{A}}_{r}}{\bm{\beta A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}}{{\partial{\theta_{j}}}}}\right]$
$\mathop{=}\limits^{\left(a\right)}\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{{Re}}}\nolimits}{\mathop{\mathrm{tr}}\nolimits}\left[\left({{\mathbf{x}}_{E}^{H}{e_{i}}e_{i}^{T}{\dot{\mathbf{A}}}_{t}^{*}{{\bm{\beta}}^{H}}{\mathbf{A}}_{r}^{H}+{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{{\bm{\beta}}^{H}}{e_{i}}e_{i}^{T}{\dot{\mathbf{A}}}_{r}^{H}}\right)\times\left({{{{\dot{\mathbf{A}}}}_{r}}{e_{j}}e_{j}^{T}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}+{{\mathbf{A}}_{r}}{\bm{\beta}}{\dot{\mathbf{A}}}_{t}^{T}{e_{j}}e_{j}^{T}{{\mathbf{x}}_{E}}}\right)\right]$
$=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{{Re}}}\nolimits}{\mathop{\mathrm{tr}}\nolimits}\left[{{\mathbf{x}}_{E}^{H}{e_{i}}e_{i}^{T}{\mathbf{\dot{A}}}_{t}^{*}{{\bm{\beta}}^{H}}{\mathbf{A}}_{r}^{H}{{{\mathbf{\dot{A}}}}_{r}}{e_{j}}e_{j}^{T}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}+{\mathbf{x}}_{E}^{H}{e_{i}}e_{i}^{T}{\mathbf{\dot{A}}}_{t}^{*}{{\bm{\beta}}^{H}}{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{\dot{A}}}_{t}^{T}{e_{j}}e_{j}^{T}{{\mathbf{x}}_{E}}}\right$.
$+\left.{{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{{\bm{\beta}}^{H}}{e_{i}}e_{i}^{T}{\mathbf{\dot{A}}}_{r}^{H}{{{\mathbf{\dot{A}}}}_{r}}{e_{j}}e_{j}^{T}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}+{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{{\bm{\beta}}^{H}}{e_{i}}e_{i}^{T}{\mathbf{\dot{A}}}_{r}^{H}{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{\dot{A}}}_{t}^{T}{e_{j}}e_{j}^{T}{{\mathbf{x}}_{E}}}\right]$
$\mathop{=}\limits^{\left(b\right)}\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{{Re}}}\nolimits}\left[{e_{i}^{T}\left({{\mathbf{A}}_{r}^{H}{{{\mathbf{\dot{A}}}}_{r}}}\right){e_{j}}e_{j}^{T}\left({{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}{\mathbf{x}}_{E}^{H}{\mathbf{\dot{A}}}_{t}^{*}{{\bm{\beta}}^{H}}}\right){e_{i}}+e_{i}^{T}\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right){e_{j}}e_{j}^{T}\left({{\bm{\beta}}{\mathbf{\dot{A}}}_{t}^{T}{{\mathbf{x}}_{E}}{\mathbf{x}}_{E}^{H}{\mathbf{\dot{A}}}_{t}^{*}{{\bm{\beta}}^{H}}}\right){e_{i}}}\right$.
$\left.{+e_{i}^{T}\left({{\mathbf{\dot{A}}}_{r}^{H}{{{\mathbf{\dot{A}}}}_{r}}}\right){e_{j}}e_{j}^{T}\left({{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{{\bm{\beta}}^{H}}}\right){e_{i}}+e_{i}^{T}\left({{\mathbf{\dot{A}}}_{r}^{H}{{\mathbf{A}}_{r}}}\right){e_{j}}e_{j}^{T}\left({{\bm{\beta}}{\mathbf{\dot{A}}}_{t}^{T}{{\mathbf{x}}_{E}}{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{{\bm{\beta}}^{H}}}\right){e_{i}}}\right]$
$\mathop{=}\limits^{\left(c\right)}\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{{Re}}}\nolimits}\left[{L{{\left({{\mathbf{A}}_{r}^{H}{{{\mathbf{\dot{A}}}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\mathbf{\dot{A}}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)}_{ij}}+L{{\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\mathbf{\dot{A}}}_{t}^{H}{{\mathbf{R}}^{*}}{{{\mathbf{\dot{A}}}}_{t}}{\bm{\beta}}}\right)}_{ij}}}\right$.
$\left.{+L{{\left({{\mathbf{\dot{A}}}_{r}^{H}{{{\mathbf{\dot{A}}}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)}_{ij}}+L{{\left({{\mathbf{\dot{A}}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{{\mathbf{\dot{A}}}}_{t}}{\bm{\beta}}}\right)}_{ij}}}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\left[{{\mathop{\mathrm{Re}}\nolimits}\left({{{\mathbf{F}}_{11}}}\right)}\right]_{ij}}$ (27)
where ${F_{11}}=\left({{\mathbf{A}}_{r}^{H}{{\dot{\mathbf{A}}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)+\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\dot{\mathbf{A}}}_{t}}{\bm{\beta}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\dot{\mathbf{A}}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\dot{\mathbf{A}}}_{t}}{\bm{\beta}}}\right)$, $\left[{\mathbf{{B}}}\right]_{ij}$} denotes the $(i,j)$ element of the matrix $\mathbf{{B}}$, $e_{i}$ denotes the $i$-th column of the identity matrix, step $\left(a\right)$ holds by
$\frac{{\partial\left({{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}}{{\partial{\theta_{i}}}}={{\dot{\mathbf{A}}}_{r}}{e_{i}}e_{i}^{T}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}+{{\mathbf{A}}_{r}}{\bm{\beta}}{\dot{\mathbf{A}}}_{t}^{T}{e_{i}}e_{i}^{T}{{\mathbf{x}}_{E}},$ (28)
step $\left(b\right)$ holds by
${\mathop{\mathrm{{tr}}}\nolimits}\left[{\left({{\mathbf{x}}_{E}^{H}{e_{i}}e_{i}^{T}{\dot{\mathbf{A}}}_{t}^{*}{{\bm{\beta}}^{H}}{\mathbf{A}}_{r}^{H}}\right)\left({{{{\dot{\mathbf{A}}}}_{r}}{e_{j}}e_{j}^{T}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}\right]$
$=e_{i}^{T}\left({{\mathbf{A}}_{r}^{H}{{{\dot{\mathbf{A}}}}_{r}}}\right){e_{j}}e_{j}^{T}\left({{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}{\mathbf{x}}_{E}^{H}{\dot{\mathbf{A}}}_{t}^{*}{{\bm{\beta}}^{H}}}\right){e_{i}}$
$=L{\left({{\mathbf{A}}_{r}^{H}{{{\dot{\mathbf{A}}}}_{r}}}\right)_{ij}}{\left({{{\bm{\beta}}^{*}}{\dot{\mathbf{A}}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)_{ij}},$ (29)
and step $\left(c\right)$ holds by
$L{\left({{\mathbf{A}}_{r}^{H}{{{\dot{\mathbf{A}}}}_{r}}}\right)_{ij}}{\left({{{\bm{\beta}}^{*}}{\dot{\mathbf{A}}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)_{ij}}$ $={\left[{{\mathbf{A}}_{r}^{H}{{{\dot{\mathbf{A}}}}_{r}}\odot{{\bm{\beta}}^{*}}{\dot{\mathbf{A}}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right]_{ij}}.$ (30)
2) The element of ${\mathbf{F}}\left({{\bm{\theta}},{{\bm{{\bm{\beta}}}}_{R}}}\right)$ is expressed as
$F\left({{\theta_{i}},{b_{{R_{j}}}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}{\mathop{\mathrm{tr}}\nolimits}\left[{\frac{{\partial{{\left({{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}^{H}}}}{{\partial{\theta_{i}}}}\frac{{\partial\left({{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}}{{\partial{b_{{R_{j}}}}}}}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}\left[L{{\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\dot{\mathbf{A}}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)}_{ij}}\right$.
$\left.+L{{\left({{\dot{\mathbf{A}}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}^{t}}}\right)}_{ij}}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\left[{{\mathop{\mathrm{Re}}\nolimits}\left({{{\mathbf{F}}_{12}}}\right)}\right]_{ij}}$ (31)
where ${F_{12}}=\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}\dot{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)+\left({\dot{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)$ and $\frac{{\partial\left({{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}}{{\partial{b_{{R_{j}}}}}}={{\mathbf{A}}_{r}}{e_{j}}e_{j}^{T}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}$.
3) The element of ${\mathbf{F}}\left({{\bm{\theta}},{{\bm{{\bm{\beta}}}}_{I}}}\right)$, ${\mathbf{F}}\left({{{\bm{{\bm{\beta}}}}_{R}},{\bm{\theta}}}\right)$}, ${\mathbf{F}}\left({{{\bm{{\bm{\beta}}}}_{R}},{{\bm{{\bm{\beta}}}}_{R}}}\right)$}, ${\mathbf{F}}\left({{{\bm{{\bm{\beta}}}}_{R}},{{\bm{{\bm{\beta}}}}_{I}}}\right)$}, ${\mathbf{F}}\left({{{\bm{{\bm{\beta}}}}_{I}},{\bm{\theta}}}\right)$}, ${\mathbf{F}}\left({{{\bm{{\bm{\beta}}}}_{I}},{{\bm{{\bm{\beta}}}}_{R}}}\right)$}, and ${\mathbf{F}}\left({{{\bm{{\bm{\beta}}}}_{I}},{{\bm{{\bm{\beta}}}}_{I}}}\right)$} are expressed as
$F\left({{\theta_{i}},{b_{{I_{j}}}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}{\mathop{\mathrm{tr}}\nolimits}\left[{\frac{{\partial{{\left({{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}^{H}}}}{{\partial{\theta_{i}}}}\frac{{\partial\left({{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}}{{\partial{b_{{I_{j}}}}}}}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}\left[L{{\left({j{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\dot{\mathbf{A}}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)}_{ij}}\right$.
$\left.+L{{\left({j{\dot{\mathbf{A}}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{{\bm{\beta}}^{*}}{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}^{t}}}\right)}_{ij}}\right]$
$=-\frac{2}{{{\sigma^{2}}}}{\left[{{\mathop{\mathrm{I}m}\nolimits}\left({{{\mathbf{F}}_{12}}}\right)}\right]_{ij}},$ (32)
$F\left({{b_{{R_{i}}}},{\theta_{j}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}\left[L{{\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{{\dot{\mathbf{A}}}}_{t}}{\bm{\beta}}}\right)}_{ij}}\right$.
$\left.+L{{\left({{\mathbf{A}}_{r}^{H}{{{\dot{\mathbf{A}}}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)}_{ij}}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\left[{{{{\mathop{\mathrm{Re}}\nolimits}}^{T}}\left({{{\mathbf{F}}_{12}}}\right)}\right]_{ij}},$ (33)
$F\left({{b_{{R_{i}}}},{b_{{R_{j}}}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}{\mathop{\mathrm{tr}}\nolimits}\left[{\left({{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{e_{i}}e_{i}^{T}{\mathbf{A}}_{r}^{H}}\right)\left({{{\mathbf{A}}_{r}}{e_{j}}e_{j}^{T}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}}\left[{L{{\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)}_{ij}}}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\left[{{\mathop{\mathrm{Re}}}\left({{{\mathbf{F}}_{22}}}\right)}\right]_{ij}},$ (34)
$F\left({{b_{{R_{i}}}},{b_{{I_{j}}}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}}{\mathop{\mathrm{tr}}}\left[{\left({{\mathbf{x}}_{E}^{H}{e_{i}}e_{i}^{T}{\mathbf{A}}_{t}^{*}{\mathbf{A}}_{r}^{H}}\right)\left({j{{\mathbf{A}}_{r}}{e_{j}}e_{j}^{T}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}}\left[{L{{\left({j{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)}_{ij}}}\right]$
$=-\frac{2}{{{\sigma^{2}}}}{\left[{{{{\mathop{\mathrm{I}m}}}^{T}}\left({{{\mathbf{F}}_{22}}}\right)}\right]_{ij}},$ (35)
$F\left({{b_{{I_{i}}}},{\theta_{j}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}\left[L{{\left({-j{\mathbf{A}}_{r}^{H}{{{\dot{\mathbf{A}}}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)}_{ij}}\right$.
$\left.+L{{\left({-j{\mathbf{A}}_{r}^{H}{{{\dot{\mathbf{A}}}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}{\bm{\beta}}}\right)}_{ij}}\right]$
$=-\frac{2}{{{\sigma^{2}}}}{\left[{{{{\mathop{\mathrm{I}m}\nolimits}}^{T}}\left({{{\mathbf{F}}_{12}}}\right)}\right]_{ij}},$ (36)
$F\left({{b_{{I_{i}}}},{b_{{R_{j}}}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}{\mathrm{tr}}\left[{\left({-j{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{e_{i}}e_{i}^{T}{\mathbf{A}}_{r}^{H}}\right)\left({{{\mathbf{A}}_{r}}{e_{j}}e_{j}^{T}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}\left[{{{L\left({-j{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)}_{ij}}}\right]$
$=-\frac{2}{{{\sigma^{2}}}}{\left[{{{{\mathop{\mathrm{I}m}\nolimits}}^{T}}\left({{{\mathbf{F}}_{22}}}\right)}\right]_{ij}},$ (37)
$F\left({{b_{{I_{i}}}},{b_{{I_{j}}}}}\right)$ $=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}{\mathrm{tr}}\left[{\left({-j{\mathbf{x}}_{E}^{H}{\mathbf{A}}_{t}^{*}{e_{i}}e_{i}^{T}{\mathbf{A}}_{r}^{H}}\right)\left({j{{\mathbf{A}}_{r}}{e_{j}}e_{j}^{T}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\mathop{\mathrm{Re}}\nolimits}\left[L{{{\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)}_{ij}}{{\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)}_{ij}}}\right]$
$=\frac{2}{{{\sigma^{2}}}}{\left[{{{{\mathop{\mathrm{Re}}\nolimits}}^{T}}\left({{{\mathbf{F}}_{22}}}\right)}\right]_{ij}},$ (38)
where ${F_{22}}=\left({{\mathbf{A}}_{r}^{H}{{\mathbf{A}}_{r}}}\right)\odot\left({{\mathbf{A}}_{t}^{H}{{\mathbf{R}}^{*}}{{\mathbf{A}}_{t}}}\right)$ and $\frac{{\partial\left({{{\mathbf{A}}_{r}}{\bm{\beta}}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}}\right)}}{{\partial{b_{{I_{j}}}}}}=j{{\mathbf{A}}_{r}}{e_{j}}e_{j}^{T}{\mathbf{A}}_{t}^{T}{{\mathbf{x}}_{E}}$.